# Boosting Generalization in Diffusion-Based Neural Combinatorial Solver via Energy-guided Sampling

### v1

@@ Abstract
Diffusion-based Neural Combinatorial Optimization (NCO) has demonstrated effectiveness in solving NP-complete (NPC) problems by learning discrete diffusion models for solution generation, eliminating hand-crafted domain knowledge. Despite their success, existing NCO methods face significant challenges in both cross-scale and cross-problem generalization, and high training costs compared to traditional solvers. While recent studies have introduced training-free guidance approaches that leverage pre-defined guidance functions for zero-shot conditional generation, such methodologies have not been extensively explored in combinatorial optimization. To bridge this gap, we propose a general energy-guided sampling framework during inference time that enhances both the cross-scale and cross-problem generalization capabilities of diffusion-based NCO solvers without requiring additional training. We provide theoretical analysis that helps understanding the cross-problem transfer capability. Our experimental results demonstrate that a diffusion solver, trained exclusively on the Traveling Salesman Problem (TSP), can achieve competitive zero-shot solution generation on TSP variants, such as Prize Collecting TSP (PCTSP) and the Orienteering Problem (OP), through energy-guided sampling across different problem scales.
@@ 1 Introduction
Figure 1: Overview of energy-guided sampling framework for achieving cross-problem generalization. Left: Pre-trained diffusion model performs denoising on original problem $G$ (TSP). Right: Proposed energy-guided sampling on target problem $G^{\prime}$ (PCTSP) conducts $K$ rewrite iterations of noise addition and guided denoising.
@@
Recent advances in deep learning have sparked interest in Neural Combinatorial Optimization (NCO), which aims to learn reusable solving strategies directly from data, eliminating the need for hand-crafted heuristics [3]. Among various deep learning approaches, diffusion-based models [4, 5] have emerged as a particularly promising direction for solving combinatorial optimization problems. These models have demonstrated remarkable capabilities in learning complex solution distributions by adapting discrete diffusion processes to graph structures. Recent works like [6, 7] have achieved state-of-the-art performance on classical problems such as the Traveling Salesman Problem (TSP), showcasing the potential of diffusion-based approaches in combinatorial optimization.
@@
However, the practical applicability of existing NCO approaches is limited by several generalization challenges. First, current models suffer from cross-scale generalization, with performance degrading significantly when applied to larger problem instances than those seen during training, especially for auto-regression-based solvers [8, 9] including transformer and reinforcement learning methods. Second, these models show limited cross-problem transfer capabilities, struggling to adapt to problem variants with modified objectives or additional constraints. While several studies have attempted to enhance learning-based solvers’ generalization through approaches such as training additional networks [10] and fine-tuning [11], these methods require substantial computational costs for training separate models for each problem type and scale.
@@
In parallel, recent advances in diffusion models, particularly in computer vision, have demonstrated the effectiveness of training-free guidance approaches for enhancing conditional generation [12, 13, 14, 15]. These approaches leverage pre-defined guidance functions or pre-trained networks to enable zero-shot conditional generation without additional training overhead. Inspired by these developments, we explore the adaptation of energy-based guidance to address the generalization challenges in combinatorial optimization.
@@
In this work, we propose an energy-guided sampling framework (Figure 1) that enhances the generalization capabilities of diffusion-based NCO solvers without requiring additional training costs. By incorporating problem-specific objectives and constraints during inference time, this approach enables zero-shot cross-problem transfer while maintaining solution feasibility. We conduct theoretical analysis that deepen our understanding on the problem transfer capability of the energy-guided sampling framework. Through experimental evaluation of the TSP-trained diffusion model on more complex variants, the Prize Collecting TSP (PCTSP) and the Orienteering Problem (OP), we empirically demonstrate its effective zero-shot transferability across problem classes of increasing complexity, while maintaining consistent performance across different problem scales. Our work represents a significant step toward more flexible and generalizable diffusion-based combinatorial optimization solvers, potentially reducing the need for problem-specific model training while maintaining competitive performance.
@@ 2 Related Works
Diffusion-based Generative Modeling. Recent advances in generative modeling have revolutionized various domains through diverse approaches, including Variational Autoencoders (VAE) [25], Generative Adversarial Networks (GAN) [26], Diffusion models [4], and GFlowNet [27]. In particular, score-based diffusion models [4, 5, 28, 29, 30] have emerged as a powerful framework operating in continuous domains.
Beyond their state-of-the-art performance in traditional generative tasks, these models have shown remarkable potential in combinatorial optimization (CO). Pioneering work by [6] established new state-of-the-art results for the Traveling Salesman Problem (TSP) by adapting discrete diffusion models [31] to graph structures. Building upon this foundation, [7] enhanced the framework’s performance through gradient search iterations during testing. [24] proposed the first diffusion-based unsupervised learning framework.
@@
A promising alternative has emerged through training-free guidance methods [12, 13, 14, 15], which are guided by pre-trained networks or loss functions. In the context of discrete diffusion models, this approach remained largely unexplored until [7] pioneered the adaptation of loss-based guidance during inference, building upon the pre-trained discrete diffusion solver framework [6]. Our work demonstrates the significant potential of energy-guided sampling in enhancing the cross-problem generalization capabilities of diffusion-based NCO solvers.
@@ 3.2 Graph-based CO Problems
Combinatorial optimization (CO) problems on graphs are fundamental to numerous real-world applications. Following recent advances [6, 7], we address these problems by formalizing graph-based CO instances as follows.
@@
We represent each problem instance as an undirected graph $G(V,E)\in\mathcal{G}$, where $V$ and $E$ denote the vertex and edge sets, respectively. This representation encompasses both vertex selection and edge selection problems, covering a broad spectrum of practical CO scenarios. For any instance $G\in\mathcal{G}$, we define a binary decision variable $\mathbf{x}\in\mathcal{X}_{\mathcal{G}}$, where $\mathcal{X}_{\mathcal{G}}=\{0,1\}^{N}$ represents the feasible solution space. The optimization objective is to find the optimal solution $\mathbf{x}^{*}$ that minimizes a problem-specific objective function $f(\cdot;G):\{0,1\}^{N}\rightarrow\mathbb{R}$:
$\mathbf{x}^{*}=\underset{\mathbf{x}\in\mathcal{X}_{\mathcal{G}}}{\text{argmin}}\;f(\mathbf{x};G),$ (1)
where the objective function decomposes into:
$f(\mathbf{x};G)=f_{\text{cost}}(\mathbf{x};G)+\beta\cdot f_{\text{valid}}(\mathbf{x};G).$ (2)
@@
Here $f_{\text{cost}}(\cdot;G)$ measures the solution quality, and $f_{\text{valid}}(\cdot;G)$ enforces problem-specific constraints through a penalty coefficient $\beta>0$. The validity function returns 00 for feasible solutions and is strictly positive for infeasible ones.
As a concrete example, consider the classical Traveling Salesman Problem (TSP): given a complete graph $G$ with edge weights, the objective is to find a minimum-weight Hamiltonian cycle. The decision variable $\mathbf{x}\in\{0,1\}^{N}$ encodes edge selections, where $f_{\text{cost}}(\cdot;G)$ measures the total tour length and $f_{\text{valid}}(\cdot;G)$ ensures the solution forms a valid Hamiltonian cycle following [6]. For the Prize Collecting TSP (PCTSP) that we considered, each vertex has a prize $r_{i}>0$ and penalty $p_{i}>0$. $f_{\text{cost}}(\cdot;G)$ measures the difference between total tour length and uncollected penalties, while $f_{\text{valid}}(\cdot;G)$ ensures the total prizes are greater than the constraint and the solution forms a valid Hamiltonian cycle. For the Orienteering Problem (OP), $f_{\text{cost}}$ aims to maximize the total collected scores from visited nodes, while $f_{\text{valid}}(\cdot;G)$ ensures both tour connectivity and the total length within the given budget.
3.3 Probabilistic Modeling for CO
@@
To leverage recent advances in deep generative models, we reformulate the CO objective through an energy-based perspective [35]. Specifically, we establish an energy function $\mathcal{E}(\cdot;G):=|y-f(\cdot;G)|$ that maps each solution to its corresponding energy state. This energy-based formulation naturally leads to a probabilistic framework through the Boltzmann distribution [36]:
$p(y|\mathbf{x};G)=\frac{\exp\big{(}-\frac{1}{\tau}\mathcal{E}(y,\mathbf{x};G)\big{)}}{\mathcal{Z}},\;\text{where}\;\mathcal{Z}=\sum_{\mathbf{x}}\exp\Big{(}-\frac{1}{\tau}\mathcal{E}(y,\mathbf{x};G)\Big{)},$ (3)
where $\tau$ controls the temperature of the system and $\mathcal{Z}$ denotes the partition function that normalizes the distribution.
@@
Recent works have demonstrated promising approaches to approximate this distribution using diffusion-based deep generative models by parameterizing a conditional distribution $p_{\theta}(\mathbf{x}|G)$ to minimize the energy function. Both supervised [6, 7] and unsupervised [24] learning paradigms have shown significant advances. Since our proposed training-free guidance mechanism is applicable to any pre-trained diffusion-based solver, we focus on the supervised learning framework in this work for ease of presentation.
@@
Given a training set $\mathcal{G}=\{G_{i}\}_{i=1}^{k}$ of i.i.d. problem instances with their optimal solutions $\mathbf{x}$ and the corresponding optimal objective values $y_{G}^{*}$, we optimize the model parameters $\theta$ by maximizing the likelihood of the optimal solutions
@@
4 Theoretical Results
4.1 Discrete Diffusion Generation Modeling
We adopt a discrete diffusion framework [31] to effectively sample optimal solutions from the learned distribution $p_{\theta}(\mathbf{x}|y^{*},G)$. In contrast to continuous diffusion models that employ Gaussian noise, our discrete formulation is particularly well-suited for graph-based combinatorial optimization problems [6, 7].
@@ 4.1 Discrete Diffusion Generation Modeling
To capture the structural properties of CO problems, we employ an anisotropic graph neural network architecture [21]. For a given instance $G$, the network learns to predict the clean data distribution $p_{\theta}(\mathbf{\widetilde{x}}_{0}|\mathbf{x}_{t},G)$. Taking TSP as an example, where $G$ encodes the 2D Euclidean coordinates of vertices, the network outputs a probability matrix $p_{\theta}(\mathbf{\widetilde{x}}_{0}|\mathbf{x}_{t},G)\in[0,1]^{N\times 2}$. This matrix parameterizes $N$ independent Bernoulli distributions, each corresponding to a binary decision variable in $\mathbf{\widetilde{x}}_{0}$. The reverse process during sampling follows:
@@ 4.2 Energy-guided Sampling for Problem Transfer
While training-free guidance has been extensively studied in computer vision [12, 13, 14, 15], its application to combinatorial optimization problems has only recently emerged [7]. We firstly extend this approach by introducing energy-based training-free guidance for new problem instances during inference, enabling flexible incorporation of additional constraints into pre-trained diffusion-based CO solvers and enhancing their cross-problem generalization capabilities.
@@
Let $\mathcal{G}^{\prime}=\{G^{\prime}_{i}\}_{i=1}^{n}$ denote a set of test instances representing variants of the original training problems, such as problems with additional constraints or multiple objectives. For a new instance $G^{\prime}$ with its optimal solution pair $(\mathbf{x},y_{G^{\prime}}^{*})$, we need to estimate the new reverse process $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},y_{G^{\prime}}^{*},G^{\prime})$ according to (9). Following the score estimation perspective of diffusion processes [5], we decompose the conditional score function at time step $t$ into two components:
@@
$\underbrace{\nabla_{\mathbf{x}_{t}}\log p_{\theta}(\mathbf{x}_{t}|y^{*}_{G^{\prime}},G^{\prime})}_{\text{posterior score}}$ $=\underbrace{\nabla_{\mathbf{x}_{t}}\log p_{\theta}(\mathbf{x}_{t}|G^{\prime})}_{\text{pre-trained prior score}}+\underbrace{\nabla_{\mathbf{x}_{t}}\log p_{t}(y^{*}_{G^{\prime}}|\mathbf{x}_{t},G^{\prime})}_{\text{energy-guided score}}.$ (10)
@@
From the Bayesian perspective, $p_{\theta}(\mathbf{x}_{t}|G^{\prime})$ can be understood as the prior, which contains knowledge of the pre-trained problems (i.e. TSP problem in our experimental settings), and $p_{t}(y^{*}_{G^{\prime}}|\mathbf{x}_{t},G^{\prime})$ corresponds to the guided likelihood that incorporates additional constraints or objectives of the variant problem. We sample from the posterior $p_{\theta}(\mathbf{x}_{t}|y^{*}_{G^{\prime}},G^{\prime})$ to generate high-quality solutions to the variant of the original training problems.
@@
Drawing upon this theoretical framework, we leverage the pre-trained diffusion model to estimate the first term $\nabla_{\mathbf{x}}\log p_{\theta}(\mathbf{x}_{t}|G^{\prime})$. In the context of cross-problem transfer, while the pre-trained model yields only a biased score function for new problem instances, we analyze their underlying connection for TSP variants in the subsequent subsection.
Meanwhile, we compute the second energy-guided term $\nabla_{\mathbf{x}_{t}}\log p_{t}(y^{*}_{G^{\prime}}|\mathbf{x}_{t},G^{\prime})$ using an energy function that specifically accounts for the additional objectives and constraints of the variant problem:
$\nabla_{\mathbf{x}_{t}}\log p_{t}(y^{*}_{G^{\prime}}|\mathbf{x}_{t},G^{\prime})\propto-\nabla_{\mathbf{x}_{t}}\mathcal{E}(y^{*}_{G^{\prime}},\mathbf{x}_{0}(\mathbf{x}_{t});G^{\prime}),$ (11)

### v2

# Boosting Generalization in Diffusion-Based Neural Combinatorial Solver via Inference Time Adaptation

@@ Abstract
Diffusion-based Neural Combinatorial Optimization (NCO) has demonstrated effectiveness in solving NP-complete (NPC) problems by learning discrete diffusion models for solution generation, eliminating hand-crafted domain knowledge. Despite their success, existing NCO methods face significant challenges in both cross-scale and cross-problem generalization, and high training costs compared to traditional solvers. While recent studies on diffusion models have introduced training-free guidance approaches that leverage pre-defined guidance functions for conditional generation, such methodologies have not been extensively explored in combinatorial optimization. To bridge this gap, we propose a training-free inference time adaptation framework (DIFU-Ada) that enables both the zero-shot cross-problem transfer and cross-scale generalization capabilities of diffusion-based NCO solvers without requiring additional training. We provide theoretical analysis that helps understanding the cross-problem transfer capability. Our experimental results demonstrate that a diffusion solver, trained exclusively on the Traveling Salesman Problem (TSP), can achieve competitive zero-shot transfer performance across different problem scales on TSP variants, such as Prize Collecting TSP (PCTSP) and the Orienteering Problem (OP), through inference time adaptation.
@@ 1 Introduction
Recent advances in deep learning have sparked interest in Neural Combinatorial Optimization (NCO), which aims to learn reusable solving strategies directly from data, eliminating the need for hand-crafted heuristics [3]. Among various deep learning approaches, diffusion-based models [4, 5] have emerged as a particularly promising direction for solving combinatorial optimization problems. These models have demonstrated remarkable capabilities in learning complex solution distributions by adapting discrete diffusion processes to graph structures [6]. Recent works like [7, 8] have achieved state-of-the-art performance on classical problems such as the Traveling Salesman Problem (TSP), showcasing the potential of diffusion-based generative approaches in combinatorial optimization.
@@
However, the practical applicability of existing NCO approaches is limited by several generalization challenges. First, current models suffer from cross-scale generalization, with performance degrading significantly when applied to larger problem instances than those seen during training, especially for auto-regression-based solvers [9, 10] including transformer and reinforcement learning methods. Second, these models show limited cross-problem transfer capabilities, struggling to adapt to problem variants with modified objectives or additional constraints. While several studies have attempted to enhance learning-based solvers’ generalization through approaches such as training additional networks [11] and fine-tuning [12], these methods require substantial computational costs and training data for training separate models for each problem type and scale.
@@
In parallel, recent advances in diffusion models, particularly in computer vision, have demonstrated the effectiveness of training-free guidance approaches for enhancing conditional generation [13, 14, 15, 16, 17]. These approaches leverage plug-and-play guidance functions or pre-trained networks to enable conditional generation without additional training overhead. These approaches demonstrate significant potential for modifying the inference phase to achieve controllable sampling from pre-trained diffusion models. Inspired by these developments, we explore the training-free inference time adaptation to address the cross-problem transfer challenges in neural combinatorial optimization in Figure 1.
Figure 1: Overview of Inference Time Adaptation for achieving zero-shot cross-problem generalization, shifting from pre-trained problem $G$ (TSP) solution space to target problem $G^{\prime}$ (PCTSP).
Figure 2: The proposed Inference Time Adaptation framework. This approach combines (1) energy-guided sampling, which incorporates problem-specific objectives and constraints, with (2) a recursive denoising-decoding travel strategy for solution refinement, enabling zero-shot cross-problem transfer without training. The Optimality Gap ($\downarrow$) on PCTSP-20 is reduced from $19.21\%$ to $4.20\%$.
@@
This work introduces an inference time adaptation framework (illustrated in Figure 2) designed to enhance the generalization capabilities of diffusion-based Neural Combinatorial Optimization (NCO) solvers without additional training costs. We scale up the inference time by combining two key components: (1) energy-guided sampling, which integrates problem-specific objective functions and constraints into the denoising process, and (2) a recursive denoising-decoding strategy applied at intermediate noise levels. This inference time adaptation framework enables zero-shot cross-problem transfer while maintaining solution feasibility. Furthermore, we present a theoretical analysis to deepen the understanding of the problem transfer capability afforded by this inference time adaptation when applied to pre-trained models. Through experimental evaluation of the TSP-trained diffusion model on more complex variants, the Prize Collecting TSP (PCTSP) and the Orienteering Problem (OP), we empirically demonstrate its effective zero-shot transferability across different problems of increasing complexity, while maintaining consistent performance across different problem scales. Our work represents a significant step toward more flexible and generalizable diffusion-based combinatorial optimization solvers, reducing the need for problem-specific model training while maintaining competitive performance.
@@ 2 Related Works
Diffusion-based Generative Modeling. Recent advances in generative modeling have revolutionized various domains through diverse approaches, including Variational Autoencoders (VAE) [27], Generative Adversarial Networks (GAN) [28], Diffusion models [4], and GFlowNet [29]. In particular, score-based diffusion models [4, 5, 30, 31, 32, 33] have emerged as a powerful framework operating in continuous domains.
Beyond their state-of-the-art performance in traditional generative tasks, these models have shown remarkable potential in combinatorial optimization (CO). Pioneering work by [6] established new state-of-the-art results for the Traveling Salesman Problem (TSP) by adapting discrete diffusion models [34] to graph structures. Building upon this foundation, [7] enhanced the framework’s performance through gradient search and the current state-of-the-art [8] introduced consistency models to achieve single-step generation. [26] proposed the first diffusion-based unsupervised learning framework.
@@
A promising alternative has emerged through training-free guidance methods [13, 14, 15, 16], which are guided by pre-trained networks or loss functions. In the context of discrete diffusion models, this approach remained largely unexplored until [7] pioneered the adaptation of loss-based guidance during inference, building upon the pre-trained discrete diffusion solver framework [6]. Our work demonstrates the significant potential of modifying the infernce phase in enhancing the zero-shot cross-problem transfer of diffusion-based NCO solvers.
@@ 3.1 Graph-based CO Problems
Combinatorial optimization (CO) problems on graphs are fundamental to numerous real-world applications. Following recent advances [6, 7], we address these problems by formalizing graph-based CO instances. Following [6, 7, 8], we represent each problem instance as an undirected graph $G(V,E)\in\mathcal{G}$, where $V$ and $E$ denote the vertex and edge sets, respectively. This representation encompasses both vertex selection and edge selection problems, covering a broad spectrum of practical CO scenarios. For any instance $G\in\mathcal{G}$, we define a binary decision variable $\mathbf{x}\in\mathcal{X}_{\mathcal{G}}$, where $\mathcal{X}_{\mathcal{G}}=\{0,1\}^{N}$ represents the feasible solution space. The optimization objective is to find the optimal solution $\mathbf{x}^{*}$ that minimizes a problem-specific objective function $\phi(\cdot;G):\{0,1\}^{N}\rightarrow\mathbb{R}$:
$\mathbf{x}^{*}=\underset{\mathbf{x}\in\mathcal{X}_{\mathcal{G}}}{\text{argmin}}\;\phi(\mathbf{x};G),$ (1)
where the objective function decomposes into a log-barrier formulation [36]:
$\phi(\mathbf{x};G)=f_{\text{cost}}(\mathbf{x};G)-\mu\sum_{i=1}^{m}\log\big{(}-g_{i}(\mathbf{x};G)\big{)},$ (2)
@@
Here $f_{\text{cost}}(\cdot;G)$ measures the solution quality, and $g_{i}(\cdot;G)$ enforces problem-specific constraints through a penalty coefficient $\mu>0$. The validity functions $g_{i}(\cdot;G)$ returns 00 for feasible solutions and is strictly positive for infeasible ones.
3.2 Probabilistic Modeling for CO Problems
@@
To leverage recent advances in deep generative models, we reformulate the CO objective through an energy-based perspective [37]. Specifically, we establish an energy function:
$\mathcal{E}(\cdot;G):=\rho(y,\phi(\cdot;G)),$ (3)
where $\rho$ represents any loss functions (cross entropy loss, L2-norm), and the energy function maps each solution to its corresponding energy state. This energy-based formulation naturally leads to a probabilistic framework through the Boltzmann distribution [38]:
$p(y|\mathbf{x};G)=\frac{\exp\big{(}-\frac{1}{\tau}\mathcal{E}(y,\mathbf{x};G)\big{)}}{\mathcal{Z}},\quad\text{where}\;\mathcal{Z}=\sum_{\mathbf{x}}\exp\Big{(}-\frac{1}{\tau}\mathcal{E}(y,\mathbf{x};G)\Big{)},$ (4)
@@
where $\tau$ controls the temperature of the system and $\mathcal{Z}$ denotes the partition function that normalizes the distribution. Recent works have demonstrated promising approaches to approximate this distribution using diffusion-based deep generative models by parameterizing a conditional distribution $p_{\theta}(\mathbf{x}|G)$ to minimize the energy function. Both supervised [6, 7, 8] and unsupervised [26] learning paradigms have shown significant advances. Since our proposed training-free guidance mechanism is applicable to any pre-trained diffusion-based solver, we focus on the supervised learning framework in this work.
@@
Given a training set $\mathcal{G}=\{G_{i}\}_{i=1}^{k}$ of i.i.d. problem instances with their optimal solutions $\mathbf{x}$ and the corresponding optimal objective values $y_{G}^{*}$, we use generative model and optimize the model parameters $\theta$ by maximizing the likelihood of the optimal solutions:
@@
3.3 Discrete Diffusion Generative Models
We adopt a discrete diffusion framework [34] to effectively sample optimal solutions from the learned distribution $p_{\theta}(\mathbf{x}|y^{*},G)$. In contrast to continuous diffusion models that employ Gaussian noise, the discrete formulation is particularly well-suited for graph-based combinatorial optimization problems [6, 7, 8].
@@ 3.3 Discrete Diffusion Generative Models
To capture the structural properties of CO problems, we employ an anisotropic graph neural network architecture [23]. For a given instance $G$, the network learns to predict the clean data distribution $p_{\theta}(\mathbf{\widetilde{x}}_{0}|\mathbf{x}_{t},G)$. Taking TSP as an example, where $G$ encodes the 2D Euclidean coordinates of vertices, the network outputs an adjacency matrix $p_{\theta}(\mathbf{\widetilde{x}}_{0}|\mathbf{x}_{t},G)\in[0,1]^{N\times 2}$. This matrix parameterizes $N$ independent Bernoulli distributions, each corresponding to a binary decision variable in $\mathbf{\widetilde{x}}_{0}$. The reverse process during sampling follows:
@@
A decoding strategy is essential to construct feasible solutions from the generated adjacency matrix, which is often represented as a heatmap from which an optimal solution is decoded. Among such strategies, greedy decoding is a straightforward approach that iteratively selects nodes or edges with the highest probabilities from this heatmap until a termination condition is met (e.g., forming a complete cycle in Traveling Salesperson Problem (TSP) applications).
4 Proposed Methods: Inference Time Adaptation
@@ 4.1 Energy-guided Sampling for Problem Transfer
While training-free guidance has been extensively studied in computer vision [13, 14, 15, 16, 17], its application to combinatorial optimization problems has only recently been explored by [7, 8]. Adopted by [7, 8], we extend this approach by introducing energy-based training-free guidance for zero-shot cross-problem transfer during sampling, enabling flexible incorporation of additional problem-specific objective functions and constraints into pre-trained diffusion-based CO solvers.
Figure 3: Visualization of the recursive denoising-decoding travel strategy. This approach applies re-noising and energy-guided denoising within an efficient intermediate stage (Time Stage 2), based on the observation that the efficacy of energy guidance varies across different noise levels.
@@
Let $\mathcal{G}^{\prime}=\{G^{\prime}_{i}\}_{i=1}^{n}$ denote a set of new problem instances which differs from the pre-trained problem set $\mathcal{G}$. For a new instance $G^{\prime}$ with its optimal solution pair $(\mathbf{x},y_{G^{\prime}}^{*})$, we need to estimate the new reverse process $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},y_{G^{\prime}}^{*},G^{\prime})$ according to (10). Following the score estimation perspective of diffusion processes [5], we decompose the conditional score function at time step $t$ into two components:
@@
$\underbrace{\nabla_{\mathbf{x}_{t}}\log p_{\theta}(\mathbf{x}|y^{*}_{G^{\prime}},G^{\prime})}_{\text{Posterior Score}}$ $=\underbrace{\nabla_{\mathbf{x}_{t}}\log p_{\theta}(\mathbf{x}_{t}|G^{\prime})}_{\text{Pre-trained Prior Score}}+\underbrace{\nabla_{\mathbf{x}_{t}}\log p_{t}(y^{*}_{G^{\prime}}|\mathbf{x}_{t},G^{\prime})}_{\text{Energy-guided Score}}.$ (11)
@@
From the Bayesian perspective, $p_{\theta}(\mathbf{x}_{t}|G^{\prime})$ can be understood as the prior, which contains knowledge of the pre-trained problems (i.e. TSP problem in our experimental settings), and $p_{t}(y^{*}_{G^{\prime}}|\mathbf{x}_{t},G^{\prime})$ corresponds to the adjustive likelihood that incorporates additional constraints or objectives of the variant problem. We sample from the posterior $p_{\theta}(\mathbf{x}_{t}|y^{*}_{G^{\prime}},G^{\prime})$ to generate high-quality solutions to the new problems.
@@
Drawing upon this theoretical framework, we leverage the pre-trained diffusion model to estimate the first term $\nabla_{\mathbf{x}}\log p_{\theta}(\mathbf{x}_{t}|G^{\prime})$. In the context of cross-problem transfer, while the pre-trained model yields only a biased score function for new problem instances, we compute the second energy-guided term $\nabla_{\mathbf{x}_{t}}\log p_{t}(y^{*}_{G^{\prime}}|\mathbf{x}_{t},G^{\prime})$ to adjust using an energy function that specifically accounts for the additional objectives and constraints of the new problems:
$\nabla_{\mathbf{x}_{t}}\log p_{t}(y^{*}_{G^{\prime}}|\mathbf{x}_{t},G^{\prime})$ $\propto-\nabla_{\mathbf{x}_{t}}\mathcal{E}(y^{*}_{G^{\prime}},\widetilde{\mathbf{x}}_{0}(\mathbf{x}_{t});G^{\prime}),$ (12)
$\propto-\nabla_{\mathbf{x}_{t}}\rho(y^{*}_{G^{\prime}},\phi(\widetilde{\mathbf{x}}_{0}(\mathbf{x}_{t});G^{\prime})),$